OpenStudy (anonymous):

f(x)=-(x+3)(x-5) finding min and max x/y intercept

OpenStudy (anonymous):

you're in calc too? lol

OpenStudy (anonymous):

right i need help

OpenStudy (anonymous):

i suck at calc lols i think u set both -x-3 and x-5 to zero and solve

OpenStudy (anonymous):

with calc you can find min/max using criticle point

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

to find criticel point , take derivative and set what you get to zero

OpenStudy (anonymous):

can u write it out..

OpenStudy (anonymous):

please

OpenStudy (anonymous):

he means critical points

OpenStudy (anonymous):

y=-(x+3)(x-5) let's use product rule to differentiate - (x+3)'(x-5)+(x-3)(x+5)'= - 1(x-5)+(x-3)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

king explains it well here

OpenStudy (anonymous):

idont udnerstand kings way. haha

OpenStudy (anonymous):

can u actaully help me with all of number u on this attatched file

OpenStudy (anonymous):

i understood it :(

OpenStudy (anonymous):

can u guys help me with wat i attached above i need help on number 6a-f

OpenStudy (anonymous):

what do you think is domain for 6) a) in other word is there any number that you can't put in for x?

OpenStudy (anonymous):

anything above 6 ?

OpenStudy (anonymous):

why not anything below 6 into f(x)=5+3x

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

could it be anything

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

why

OpenStudy (anonymous):

you can put anything number and you won't get in problem so your domain is all real number

OpenStudy (anonymous):

wwhat about range

OpenStudy (anonymous):

domain is number you put in for x , range is number you get back as y as result of you putting value for x

OpenStudy (anonymous):

if you can put all value for x, you will get any value of y so range is also all real number

OpenStudy (anonymous):

ok now b

OpenStudy (anonymous):

b is same thing

OpenStudy (anonymous):

so is c

OpenStudy (anonymous):

d) is different tell me what number can't we put in?

OpenStudy (anonymous):

error

OpenStudy (anonymous):

shows up

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

in the calculator ..

OpenStudy (anonymous):

for what number did you get an error?

OpenStudy (anonymous):

-1 -2 -3 -4 and all negative

OpenStudy (anonymous):

ok , so we can't put negative number in there , 0 is fine so domain is x must be greater or equal to 0 x>=0

OpenStudy (anonymous):

how bout range

OpenStudy (anonymous):

square root only return positive values sqrt[0]=0 sqrt[1]=1 sqrt[4]=2 ect....... so range is too y>=0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

alright, I forgot to do range for b and c

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

y= x^2 y of that function goes from 0 to infinity if we shift that up y=x^2+4 y goes from 4 to infinity

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for c) well just look the graph http://www.wolframalpha.com/input/?i=(1.5)%5Ex&t=crmtb01 goes from 0 to infinity

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

alright.

OpenStudy (anonymous):

i need e nd f

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

any number for domain

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

range?

OpenStudy (anonymous):

y=-x^2 goes from -negative infinity to 0 so y=-x^2+2 is shifted up by two

OpenStudy (anonymous):

how about f

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what numbers doesn't works?

OpenStudy (anonymous):

any number below o

OpenStudy (anonymous):

wait can u help me to explain number 5 i got the answers but idk how to explain the meaning

OpenStudy (anonymous):

no, only number that doesn't work is 0

OpenStudy (anonymous):

how about range

OpenStudy (anonymous):

all real number

OpenStudy (anonymous):

i need help explainging what 5 means a-d

OpenStudy (anonymous):

make another thread